OpenStudy (anonymous):

Find the quotient of the complex numbers. If necessary, use the slash bar ( / ) to enter a fraction.(2 + 2i )/(5 + 4i )

OpenStudy (apoorvk):

Same thing once again! Just multiply the numerator and denominator once agin with the conjugate of the denominator! and simplify!

OpenStudy (apoorvk):

*again

OpenStudy (anonymous):

and don't think too hard about the denominator, because it is always \(a^2+b^2\) in your case \(5^2+4^2\)

OpenStudy (anonymous):

=(2+2i)/(5+4i) * (5-4i)/(5-4i) =(2+2i)(5-4i) / (5)^2-(4i)^2 =2(5-4i)+2i(5-4i) / 25-16i^2 =10-8i+10i-8i^2 / 25-(-16) =10+2i-8(-1) / 25+16 =18+2i/41